OpenStudy (anonymous):

Is there a fast way to factor 4x^2-x+18?

OpenStudy (anonymous):

sorry is -18 not +18

OpenStudy (anonymous):

what i do ix 4 x 18 and then try out all the multiples but i remember there is a cross method way? i'm not sure

OpenStudy (anonymous):

(4x-9)(x+2)

OpenStudy (anonymous):

i know the answer..asking how to solve it in a faster way

OpenStudy (anonymous):

i like ac method - pretty quick for all those not in the special squares categories for me

Directrix (directrix):

I learned the "bust the b" technique. I can't say that it is fast. 4 x² -1x + 18 a = 4 and c = 18 and b = -1. The task on "bust the b" is to find numbers that mutiply to a*c AND add to b. In this case, numbers that multiply to 4*18 and add to -1. 14 * 18 breaks aparts into the product of prime factors : 2*2*2*3*3 Look at those five factor of 14 * 18 and try to put them in the form of 2 numbers that differ by -1. 2*2*2 and 3*3 appear to differ by 1. That would be 8 and 9. "b" has been "busted into 8 and - 9 whose sum is -1 which equals "b" in this problem. 4x - x+ 18 = 4 x² + 8 x - 9 x - 1. The "b" has been busted. Factoring by grouping comes next. 4 x² + 8 x - 9 x - 18 = 4x ( x + 2) - 9 ( x + 2 ) = --> On this step, (x + 2) is the common factor of the expression. [ (x + 2) ] ( 4x - 9 ) = ( x + 2 ) ( 4x - 9)